As part of a lawsuit settlement, a major corporation offers you $75,000 today or $100,000 next year. Which do you choose if interest rates are 5 percent? If they are 15 percent?
Answer #1

Present value of both the amounts should be compared to take decision

If the interest rate is 5%,

Present value of amount received today =$75,000

Next year = 100,000/1.05 =$95,238.10

Hence, choose 100,000 next year

If rate =15%

Today =$75,000

Next year = 100,000/1.15 =$86,956.52

Hence, choose 100,000 next year